40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 17 of 335  Not logged in ELOG logo
ID Date Author Type Category Subject
  16052   Mon Apr 19 21:54:55 2021 YehonathanUpdateCDSUpdated c1auxey wiring plan

Except for the feed-throughs that require a DB9-M connector I finished wiring and labeling the Acromag, following Jon's updated wiring plan.

We can start testing the differential inputs until the missing connectors arrive.

 

  16051   Mon Apr 19 19:40:54 2021 YehonathanUpdateGeneralGlue Freezer completely frozen

{Paco, Yehonathan}

We broke the last chunk of ice and cleaned the fridge. We move the fridge back inside and plugged it into the wall. The glues were moved back from the main fridge.

The batteries that were found soaking wet are now somewhat dry and were left on the cabinet drawers for future recycling.

Quote:

{Anchal, Paco, Yehonathan}

We took the glue fridge outside.

 

  16050   Mon Apr 19 13:15:20 2021 Ian MacMillanUpdateCDS40m LSC simPlant model

The x1SUSsim model on the docker was made in a more recent version of Simulink so I updated Matlab (see this)

I updated Matlab to 2021a so now the docker has 2020b and 2021a installed. This should also install Simulink 10.3 for the sus model to open. I used my account to activate it but I can change it over if I get a dedicated license for this. I am not sure what Jon did for the 2020b that he installed.

it is giving me "License error: -9,57" so I guess it didn't work... I will try to just make the model on the 2020b just so I have something.

-----------------------------------------------------------------------------------------------------------------------------

I was able to fix the problem with the activation (using this).

I can now open the x1sussim.mdl file. It is trying to access a BSFM_MASTER Library that it doesn't seem to have. the other files don't seem to have any warnings though. 

the simple suspension model can be found in home/controls/docker-cymac/userapps/c1simpsus.mdl on the docker system. This is where I will put my model. (right now it is just a copied file)

Also using Simulink on the docker is very slow. I think this is either a limit of the x2goclient software or the hardware that the docker is running on.

  16049   Mon Apr 19 12:18:19 2021 Anchal, PacoUpdateSUSTested proposed filters for POS colum in MC2 output matrix

The filters were somewhat successful, how much we can see in attachment 1. The tip about difference between eigenmode basis and cartesian basis was the main thing that helped us take data properly. We still used OSEM data but rotated the output from POS, PIT, YAW to x, theta, phi (cartesian basis where x is also measured as angle projected by suspension length).


Eigenmode basis and Cartesian basis:

  • It is important to understand the difference between these two and what channels/sensors read what.
  • Eigenmode basis as the name suggests is the natural basis for the suspended pendulum.
    • It signifies the motion along three independent and orthogonal modes of motion: POS (longitudinal pendulum oscillation), PIT, and YAW.
    • The position of optic can be written in eigenmode basis as three numbers:
      • POS: Angle made by the center of mass of optic with verticle line from suspension point.
      • PIT: Angle made by the optic face with the suspension wires (this is important to note).
      • YAW: Angle made by optic surface with the nominal plane of suspension wires. (the yaw angle basically).
  • Cartesian basis is the lab reference frame.
    • Here we define three variables that can also represent an optic positioned and orientation:
      • x: Angle made by the center of mass of optic with verticle line from suspension point. (Same as POS)
      • \large \theta: Angle made by the optic surface with absolute verticle (z-axis) in lab frame.
      • \large \phi: Twist of the optic around the z-axis. Same as YAW angle above.
  • We want to apply the feedback gains and filters in eigenmode basis because they are a set of known independent modes. (RXA: NOOO!!!!!! read me elog entry on this topic)
  • Hence, the output from input matrix of suspensions comes out at POS, PIT and YAW in the eigenmode basis.
  • However, the sensors of optic positional, and orientation such at MC_F, wave front sensors and optical levers measure it in lab frame and thus in cartesian basis.
  • Essentially, the \large \theta measured by these sensors is different from the PIT calculated using the OSEM sensor data and is related by:
    • \large \theta = PIT - POS, where PIT and POS both are in radians as defined above.
  • When we optimized the cross-coupling in output matrix at high frequencies using the MC_F and WFS data, we actually optimized it In cartesian basis.
  • The three feedback filters from POS, PIT and YAW which carry data in the eigenmode basis need to be rotated into the cartesian basis in the output matrix before application to the coils.
  • The so-called F2A and A2L filters are essentially doing this rotation.
  • Above the resonant frequencies, the PIT and \large \theta become identical. Hence we want our filters to go to unity

The two filter sets:

  • The filters are named Eg2Ctv1 and Eg2Ctv2 on the POS column of MC2 output matrix.
  • This is to signify that these filters convert the POS, PIT, and YAW basis data (eigenmode basis data) into the cartesian basis (x, theta, phi) in which the output matrix is already optimized at higher frequencies.
  • v1 filter used an ideal output matrix during the calculation of filter as described in 16042 (script at scripts/SUS/OutMatCalc/coilBalanceDC.py).
  • Attachment 2 shows these filter transfer functions.
  • v2 filter use the output matrix optimized to reduce cross-coupling amount cartesian basis modes (MC_F, WFS_PIT and WFS_YAW) in 16009.
  • Attachment 3 shows these filter transfer funcitons.
  • Because of this, the v2 filter is different among right and left coils as well. We do see in Attachment 1 that this version of filter helps in reducing POS->YAW coupling too.

Test procedure:

  • We uploaded both the diagonalized input matrix and the diagonalized output matrix as calculated earlier.
  • We measured channels C1:SUS-MC2_SUSPOS_IN1_DQ, C1:SUS-MC2_SUSPIT_IN1_DQ, and C1:SUS-MC2_SUSYAW_IN1_DQ throughout this test.
  • These channels give output in an eigenmode basis (POS, PIT, and YAW) and the rows of the input matrix have some arbitrary normalization.
  • We normalize these channels to have same input matrix normalization as would be for ideal matrix (2 in each row).
  • Then, assuming the UL_SENS, UR_SENS, LR_SENS, and LL_SENS channels that come at input of the input matrix are calibrated in units of um, we calculate the cartesian angles x, theta, phi. for this calculation, we used the distance between coils as 49.4 mm (got it from Koji) and length of suspension as 0.2489 m and offset of suspension points from COM, b = 0.9 mm.
  • Now that we have true measures of angles in cartesian basis, we can use them to understand the effect on cross coupling from the filters we used.
  • PSL shutter is closed and autolocker is disabled. During all data measurements, we switched of suspension damping loops. This would ensure that our low frequency excitation survives for measurement at the measurement channels.
  • We first took reference data with no excitation and no filters for getting a baseline on each channel (dotted curves in Attachment 1).
  • We then send excitation of 0.03 Hz with 500 counts amplitude at C1:SUS-MC2_LSC_EXC and switched on LSC output.
  • One set of data is taken with no filters active (dashed curve in attachment 1).
  • Then two sets of data are taken with the two filters. Each data set was of 500s in length.
  • Welch function is used to take the PSD of data with bin widht of  0.01Hz and 9 averages.

Results:

  • Filter v1 was the most successful in reducing \large x \rightarrow \theta coupling by factor of 17.5.
    • The reduction in \large x \rightarrow \phi coupling was less. By a factor of 1.4.
  • Filter v2 was worse but still did a reduction of \large x \rightarrow \theta coupling by factor of 7.8.
    • The reduction in \large x \rightarrow \phi coupling was better. By a factor of 3.3.

Next, filters in PIT columns too

  • We do have filters calculated for PIT as well.
  • Now that we know how to test these properly, we can test them tomorrow fairly quickly.
  • For the YAW column though, the filters would probably just undo the output matrix optimization as they are derived from ideal transfer function models and ideally there is no coupling between YAW and other DOFs. So maybe, we should skip putting these on.
Attachment 1: CrossCoupleTestForEgToCtFilters.pdf
CrossCoupleTestForEgToCtFilters.pdf
Attachment 2: uncFilters.pdf
uncFilters.pdf uncFilters.pdf uncFilters.pdf
Attachment 3: uncFilters_v2.pdf
uncFilters_v2.pdf uncFilters_v2.pdf uncFilters_v2.pdf
  16048   Mon Apr 19 10:52:27 2021 YehonathanUpdateGeneralGlue Freezer completely frozen

{Anchal, Paco, Yehonathan}

We took the glue fridge outside.

  16047   Mon Apr 19 09:17:51 2021 JordanUpdateVACEmpty N2 Tanks

When I came into the lab this morning, I noticed that both N2 tanks were empty. I had swapped one on Friday (4-16-21) before I left the lab. Looking at the logs, the right tank (T2) sprung a leak shortly shortly after install. I leak checked the tank coupling after install but did not see a leak. There could a leak further down the line, possibly at the pressure transducer.

The left tank (T1) emptied normally over the weekend, and I quickly swapped the left tank for a full one, and is curently at ~2700 psi. It was my understanding that if both tanks emptied, V1 would close automatically and a mailer would be sent out to the 40m group. I did not receive an email over the weekend, and I checked the Vac status just now and V1 was still open.

I will keep an eye on the tank pressure throughout the day, and will try to leak check the T2 line this afternoon, but someone should check the vacuum interlocks and verify.

 

Attachment 1: N2_Pressure.PNG
N2_Pressure.PNG
  16046   Sun Apr 18 21:29:55 2021 ranaUpdatePSLLaser amplifier
  • Ideally, we put the chiller outside of the interferometer area. The PSL chiller used to be in the control room near the door by IMC REFL. We could also put it in the drill press room.
  • Once we figure out a couple of places where the Diode Box can go, we can ask facilities to make the appropriate power connections. They will have to eval the situation to figure out if the main power to the lab needs to be shut down.
  • Can we put the laser diode box in the drill press room too? Then the hoses can be short. Perhaps less EMI getting into our sensitive places.
  16045   Fri Apr 16 19:07:31 2021 YehonathanUpdateGeneralGlue Freezer completely frozen

There is still a huge chunk of unmelted ice in the fridge. I moved the content of that fridge in the main fridge and put "do not eat" warning signs.

I returned the fridge to the lab and plugged it back in to prevent flooding.

Defrosting will have to continue on Monday.

  16044   Fri Apr 16 18:21:36 2021 YehonathanUpdatePSLLaser amplifier

I surveyed a bit the 1X1/2 area to plan for the installation of the laser amplifier.

There is a vacancy at the bottom of 1X2 (attachment 1). I measured the dimensions of the diode box (DB) and it should fit. The optical fiber bundle is 75m long and should reach the amplifier head on the table easily.

According to the specs, the maximum power consumption of the DB is 800W (typically 600W), it should probably have its own circuit breaker. It can easily draw more than a few amps. The rack power strips are connected to this 4 socket box (attachment 2), is this just another power strip? It is connected to a circuit breaker with a 30A rating. How do we proceed from here?

In any case, we will need at least 2 meters of power cable.

I also tried to find a suitable place for a water chiller. A few suggestions are in the attachments. Basically either between the electronics shelves and the small rack next to 1X2 or next to the small rack close to the optical table. Maybe put it where the ladder sits and find another place for the ladder. Other options?

We would also need a windows machine running the Beckhoff software. The idea is that all the different laser components (DB, chillers, interlocks, switches) are connected to the EtherCat (over the ethernet infrastructure) so that the Beckhoff code can recognize a failure and switch off everything.

The things that are monitored:

1. Is the NPRO on?

2. Is the flow rate from the chillers enough?

3. Is the temperature of the diodes in the normal range?

4. Is one of the interlocks open?

5. Was one of the emergency buttons pushed?

6. Was the key switch on the DB turned to OFF?

The DB is EtherCat ready but the rest of the signals need to be interfaced somehow. Do we have to buy these EtherCAT terminals?

 

 

Attachment 1: 20210416_143642.jpg
20210416_143642.jpg
Attachment 2: 20210416_145408.jpg
20210416_145408.jpg
Attachment 3: 20210416_145448.jpg
20210416_145448.jpg
Attachment 4: 20210416_181324.jpg
20210416_181324.jpg
  16043   Fri Apr 16 15:47:58 2021 ranaUpdateSUSTested proposed filters for POS colum in MC2 output matrix

Looks mostly right, but you used the OSEM sensors as readbacks. We are diagonalizing using the cavity sensors. Using the diagonalized input matrix is also good since that will reduce the cross-coupling due to the damping loops.

Its sort of a subtle issue:

  1. The sensors are are diagonalized into the eigenmode basis, not the Cartesian basis of the mirror motion.
  2. What the cavity cares about is the Cartesian motion.
  3. Q1: in the model, how much Cartesian pitch motion is there at the POS eigenfrequency in the free-swinging case?
  4. Q2: should we somehow diag the input matrix into the Cartesian basis?
  5. Q3: if so, How?
  6. Q4: and why?
  16042   Fri Apr 16 11:36:36 2021 Anchal, PacoUpdateSUSTested proposed filters for POS colum in MC2 output matrix

We tried two sets of filters on the output matrix POS column in MC2. Both versions failed. Following are some details.


How test was done:

  • PSL shutter was closed and autolocker was switch off.
  • Turned off damping on POS, PIT, and YAW using C1:SUS-MC2_SUSPOS_SW2, C1:SUS-MC2_SUSPIT_SW2, and C1:SUS-MC2_SUSYAW_SW2.
  • Reference data was taken with no excitation to get relative increase at excitation.
  • Channels C1:SUS-MC2_SUSPIT_IN1, C1:SUS-MC2_SUSPOS_IN1, and C1:SUS-MC2_SUSYAW_IN1.
  • Frist we sent an excitation through LOCKIN1 at 0.11 Hz and 500 counts amplitude.
  • LOCKIN column in MC2 output matrix was kept identical to POS column, so all ones.
  • This formed our reference data set when no filters were used. Attachment 1.
  • Note that the peak at 0.03 Hz is due to LOCKIN2 that was left switched on due to autolocker.
  • Then the calculated filters were loaded using foton. Procedure:
    • Right click on filter bank med. Got to Execute-> Foton.
    • Go to File and uncheck 'Read Only'.
    • Find the filter module name in Module drop down.
    • Select an empty module section in Sections.
    • Write a name for the filter. We used DCcoupF2A and DCcouF2A2 for the two version respectively.
    • Paste the zpk foton format in Command.
    • Check with Bode plot if these are correct filters. Then click on Save. It will take about 30s to become responsive again.
    • GO back to filter bank medm screen and click on 'Load Coefficients'. This should start displaying your new filter module.
    • To switch on the module, click on the button below its name.
  • Once fitlers were loaded, we realized we can not use the LOCKIn to excite anymore as it comes as separate excitation.
  • So we used awggui to excite C1:SUS-MC2_LSCEXC at 0.11 Hz and 500 counts.
  • Then we retook the data and checked if the peaks are visible on PIT and YAW channels and how high they are.

Filer version 1

  • This was calculated by starting from ideal output matrix elements as they are currently loaded. All 1's for POS and so on.
  • The calculations were done in scripts/SUS/OutMatCalc/coilBalanceDC.py.
  • This file uses a state space model of the suspension and calculated the cross-coupling. Then the cross coupling is inverted and applied to the current output matrix elements to get correction DC gains.
  • These corrected DC gains are then used to create the filters as described in last post.
  • Attachment 2 shows the filter transfer functions and Attachment 3 shows the test results. Failed :(.
  • There was practivally no change in cross coupling that we can see.

Filter version 2:

  • In this version we used the output matrix optimized at high frequencies earlier (16009).
  • While testing this version, we also uploaded this optimised output amtrix at high frequency.
  • In this test, we realized the LOCKIN2 was on and switched it off manually. All excitations were done through awggui.
  • Attachment 4 shows the filter transfer functions and Attachment 5 shows the test results. Failed :(.
  • There was again practivally no change in cross coupling that we can see.

Forgot to upload new MC2 input matrix:

  • In hindsight, we should have uploaded our diagonalized suspension input matrix in MC2.
  • Without it, there was cross-coupling the in the sensor data to begin with.
  • But this can only be part of the reason why all our filters failed miserably.
  • Because the output matrix was not diagonalized earlier but it was not so bad. Onyl a fresh test can tell if it was the culprit.
Attachment 1: 20210416_MC2DCcoilBalancingNoFilters.pdf
20210416_MC2DCcoilBalancingNoFilters.pdf
Attachment 2: uncFilters.pdf
uncFilters.pdf uncFilters.pdf uncFilters.pdf
Attachment 3: 20210416_MC2DCcoilBalancingWithFilters.pdf
20210416_MC2DCcoilBalancingWithFilters.pdf
Attachment 4: uncFilters_v2.pdf
uncFilters_v2.pdf uncFilters_v2.pdf uncFilters_v2.pdf
Attachment 5: 20210416_MC2DCcoilBalancingWithFilters_v2.pdf
20210416_MC2DCcoilBalancingWithFilters_v2.pdf
  16041   Fri Apr 16 11:31:00 2021 ranaUpdateelogelog stuck ~10 AM today

found it unresponsive. Restarted fine using procedure documented in wiki

  16040   Fri Apr 16 10:58:16 2021 YehonathanUpdateGeneralGlue Freezer completely frozen

{Paco, Anchal, Yehonathan}

We emptied the fridge and moved the amplifier equipment on top of the amplifier crate. We unplugged the freezer and moved it out of the lab to defrost (attachment).

Quote:

I was looking at the laser head/amp and somehow decided to open the glue freezer. And it was stuck. I've managed to open it but the upper room was completely frozen.
Some of the batteries were embedded in a block of ice. I think we should throw them out.

Can the person who comes in the morning work on defrosting?

- Coordinate with Yehonathan and move the amps and the wooden crate so that you can move the freezer.

- Remove the contents to somewhere (it's OK to be room temp for a while)

- Unplug the freezer

- Leave the freezer outside with the door open. After a while, the ice will fall without care.

- At the end of the day, move it back to the lab. Continue defrosting the other day if the ice remains.

 

Attachment 1: 20210416_105048.jpg
20210416_105048.jpg
  16039   Fri Apr 16 00:21:52 2021 KojiUpdateGeneralGlue Freezer completely frozen

I was looking at the laser head/amp and somehow decided to open the glue freezer. And it was stuck. I've managed to open it but the upper room was completely frozen.
Some of the batteries were embedded in a block of ice. I think we should throw them out.

Can the person who comes in the morning work on defrosting?

- Coordinate with Yehonathan and move the amps and the wooden crate so that you can move the freezer.

- Remove the contents to somewhere (it's OK to be room temp for a while)

- Unplug the freezer

- Leave the freezer outside with the door open. After a while, the ice will fall without care.

- At the end of the day, move it back to the lab. Continue defrosting the other day if the ice remains.

Attachment 1: P_20210416_000906.jpg
P_20210416_000906.jpg
Attachment 2: P_20210416_000850.jpg
P_20210416_000850.jpg
  16038   Thu Apr 15 18:33:39 2021 KojiUpdateCDSMore 16bit ADC adapter boards

We received 10x 16bit ADC adapter boards from Todd. S2100687~S2100696

The number of soldered resistors seems to be less than that on the schematics. They are related to duotone, so check if it's OK upon use.

Attachment 1: P_20210415_183139_1.jpg
P_20210415_183139_1.jpg
  16037   Thu Apr 15 17:24:08 2021 JonUpdateCDSUpdated c1auxey wiring plan

I've updated the c1auxey wiring plan for compatibility with the new suspension electronics. Specifically it is based on wiring schematics for the new HAM-A coil driver (D1100117), satellite amplifier (D1002818), and HV bias driver (D1900163).

Changes:

  • The PDMon, VMon, CoilEnable, and BiasAdj channels all move from DB37 to various DB9 breakout boards.
  • The DB9 cables (x2) connecting the CoilEnable channels to the coil drivers must be spliced with the dewhitening switching signals from the RTS.
  • As suggested, I added five new BI channels to monitor the state of the CoilEnable switches. For lack of a better name, they follow the naming convention C1:SUS-ETMY_xx_ENABLEMon.

@Yehonathan can proceed with wiring the chassis.

Quote:

I finished prewiring the new c1auxey Acromag chassis (see attached pictures). I connected all grounds to the DIN rail to save some wiring. The power switches and LEDs work as expected.

I configured the DAQ modules using the old windows machine. I configured the gateway to be 192.168.114.1. The host machine still needs to be setup.

Next, the feedthroughs need to be wired and the channels need to be bench tested.

Attachment 1: C1AUXEY_Chassis_Feedthroughs_-_By_Connector.pdf
C1AUXEY_Chassis_Feedthroughs_-_By_Connector.pdf
  16036   Thu Apr 15 15:54:46 2021 gautamUpdateIOOWaveplate commissioning - hardware installed

[jordan, gautam]

We did the following this afternoon.

  1. Disconnected the cable from the unused (and possibly not working) RefCav heater power supply, and removed said PS from 1X1. There was insufficient space to install the ESP300 controller elsewhere. I have stored the power supply along the east arm under the beamtube, approximately directly opposite the RFPD cabinet.
  2. Installed the ESP 300 - conveniently, the HP DCPS was already sitting on some rails and so we didn't need to add any.
  3. Ran a long D25-D25 cable from the ESP300 to the NE corner area of the PSL enclosure. The ends of the cable are labelled as "ESP end" and "Waveplate end". The HEPA was turned on for the duration we had the enclosure open, and I have now turned it off.
  4. Connected the waveplate to this cable. Also re-connected the ESP300 to the c1psl supermicro host via the USB-RS232 adapter cable.

The IMC stayed locked throughout our work, and judging by the CDS overview screen, we don't seem to have done any lasting damage, but I will run more tests. Note that the waveplate isn't yet installed in the beam path - I may do this later today evening depending on lab activity, but for now, it is just sitting on the lower shelf inside the PSL enclosure. I will post some photos later.

Quote:
 

So this system is ready to be installed once Jordan and I find some time to lay out cabling + install the ESP300 controller in a rack.


Update: The waveplate was installed. I gave it a couple of rounds of cleaning by first contact, and visually, it looked good to me. More photos uploaded. I also made some minor improvements to the MEDM screen, and setup the communication script with the ESP300 to run as a systemd service on c1psl. Let's see how stable things are... I think the philosophy at the sites is to calibrate the waveplate rotation angle in terms of power units, but i'm not sure how the unit we have performs in terms of backlash error. We can do a trial by requesting ~100 "random" angles, monitoring the power in s- and p-polatizations, and then quanitfying the error between requested and realized angles, but I haven't done this yet. I also haven't added these channels to the set recorded to frames / to the burt snapshot - do we want to record these channels long term?

  16035   Thu Apr 15 11:41:43 2021 AnchalUpdateSUSProposed filters for output matrix aka F2A aka F2P

Here' s aquick update before we leave for lunch. We have managed to calculate some filter that would go on the POS column in MC2 output matrix filter banks aka F2A aka F2P filters. In the afternoon if we can come and work on the IMC, we'll try to load them on the output matrix. We have never done that so it might take some time for us to understand on how to do that. Attached is the bode plot for these proposed filters. Let us know if you have any comments.

Attachment 1: MC2propPOSfb.pdf
MC2propPOSfb.pdf
  16034   Thu Apr 15 09:46:24 2021 YehonathanUpdatePSLLaser amplifier

Some more relevant documents provided by Matt:

Phase III:70W amplifier integration at LIGO

70W amplifier External Shutter

aLIGO PSL high power attenuator

 

  16033   Wed Apr 14 23:55:34 2021 gautamUpdateElectronicsHV Coil driver assembly

I've occcupied the southernmost electronics bench for assembling the 4 production version HV coil driver chassis. I estimate it will take me 3 days, and have left a sign indicating as much. Once the chassis assembly is done, I will need to occupy the northernmost bench where bench supplies are to run some functionality tests / noise measurements, and so unless there are objections, I will move the Acromag box which has been sitting there.

  16032   Wed Apr 14 19:48:18 2021 gautamUpdatePSLLaser amplifier

A couple of years ago, I got some info about the amplifier setup at the sites from Terra - sharing here in case there is some useful info in there (our setup will be rather different, but it looked to me like our Amp is a 2017 vintage and it may be that the performance is not the same as reported in the 2019 paper).

collection of docs (table layout in 'Proposed....setup') : https://dcc.ligo.org/LIGO-T1700046

LVC 70W presentation: https://dcc.ligo.org/LIGO-G1800538 

I guess we should double check that the beam size everywhere (in vacuum and on the PSL table) is such that we don't exceed any damage thresholds for the mirrors used. 

  16031   Wed Apr 14 17:53:38 2021 AnchalUpdateSUSPlan for calculating filter banks for output matrix aka F2A aka F2P

Plan of action

  • Get the transfer functions of the suspension plant from actuated DOF to sensed DOF. We'll verify Bhavini's state-space model and get these transfer functions. Use the model TFs, not measured.
  • For each of POS->POS, PIT->PIT, and YAW->YAW, we'll get the resonant frequency and Q of the resonance from these models. No, forget about the Q.
  • We can correct the resonant frequencies from the measured ones in our free swinging data.
  • Now, we'll repeat the following for each column of output matrix filters (inspired from scripts/SUS/F2Pcalc.py, but not fully understood how/why):
    • Select col (eg. POS)
    • Set f0 to the resonant frequency.
    • Calculate \large f_{UL} = f_0 * \sqrt{G_{UL}} where GUL is the corrected DC gain we got after output matrix optimization earlier. (Not sure how, why?). No, use the SS model.
    • Calculate fUR, fLL, and fLR like above.
    • Set \large Q_{UL} = \sqrt{G_{UL}}   (This just seems like a way of keeping some approximately low Q, ideally we should keep this same to what we got above but that might cause saturation issues like Rana mentioned in the meeting)
    • Then, set the following filter in the output matrix element for UL:
                            G_{UL}\frac{1 + i\frac{f}{f_{UL}Q_{UL}} - \frac{f^2}{f_{UL}^2}}{1 + i\frac{f}{f_{0}} - \frac{f^2}{f_{0}^2}}
      which is in zpk form equivalent to:
                            z: \frac{f_0}{2 Q_{UL}} +/- i f_0 \sqrt{1 - \frac{1}{4Q_{UL}}} \quad, \quad p: \frac{f_0}{2} +/- i f_0 \frac{\sqrt{3}}{2} \quad, \quad k: G_{UL}
    • Repeat the above for UR, LL, LR.
    • Note that this filter function takes values GUL at DC and at high frequencies while it would dip at the resonant frequency for POS with depth and narrowness directly proportional to QUL. No, the DC gain is different from the AC gain.
  • However, the F2P filter plots we found in several places on elog look a bit different. Like here: 40m/4719. One important difference is that the filter magnitude always become 1 after the resonance at higher frequencies. Yes, this is  what we want, since you already did the balancing at high frequencies.
  • A preliminary plot of the above calculation for the 1,1 output matrix filter bank (POS -> UL) is attached in Attachment 1.

Discussion:

  • We can make 12 such filters for the 12 numbers we got for the optimized output matrix. Is that the aim or should we do it only for the POS column as has been done in past?
  • We are not sure how the choice of Q is made in setting the above filter function. We'll think more about it to understand this.
  • We are also not sure how the choice of fUL is made above. It looks like depending on the correction gain, we want to slide the zero positions with respect to the pole positions which are fixed at the resonant frequency as expected. This seems to have some complex explanation.
  • Please let us know if we are planning this right before we dive into these calculations/script writing. Thanks.

Edit Thu Apr 15 08:32:58 2021 :

Comments are from Rana.

Corrected the plot in the attachment. It shows the correct behavior at high frequencies now.

Attachment 1: MC2propF2A_UL.pdf
MC2propF2A_UL.pdf
  16030   Wed Apr 14 16:46:24 2021 AnchalUpdateGeneralIFO State

That makes sense. I assumed that IFO-STATE is configured as you have proposed it to be configured. This could be implemented in later.

Quote:
 

a better way would be to configure the EPICS record to automatically set / unset itself based on some diagnostic channels. For example, the "PMC locked" bit should be set if (i) the PMC REFL is < 0.1 AND (ii) PMC TRANS is >0.65 (the exact thresholds are up for debate). Then we are truly recording the state of the IFO and not relying on some script to write to the bit (I haven't thoguht through if there are some edge cases where we need an unreasonable number of diagnostic channels to determine if we are in a certain state or not).

 

  16029   Wed Apr 14 15:30:29 2021 ranaUpdateGeneralSorry, it was me

Maybe tighten the tensioner on the door closer so that it closes by itself even in the low velocity case. Or maybe just use the front door like everyone else?

  16028   Wed Apr 14 14:52:42 2021 gautamUpdateGeneralIFO State

The C1:IFO-STATE variable is actually a bunch (16 to be precise) of bits, and the byte they form (2 bytes) converted to decimal is what is written to the EPICS channel. It was reported on the call today that the nominal value of the variable when the IMC is locked was "8", while it has become "10" today. In fact, this has nothing to do with the IMC. You can see that the "PMC locked" bit is set in Attachment #1. This is done in the AutoLock.sh PMC autolocker script, which was run a few days ago. Nominally, I just lock the PMC by moving some sliders, and I neglect to set/unset this bit.

Basically, there is no anomalous behavior. This is not to say that the situation cannot be improved. Indeed, we should get rid of the obsolete states (e.g. FSS Locked, MZ locked), and add some other states like "PRMI locked". While there is nothing wrong with setting these bits at the end of execution of some script, a better way would be to configure the EPICS record to automatically set / unset itself based on some diagnostic channels. For example, the "PMC locked" bit should be set if (i) the PMC REFL is < 0.1 AND (ii) PMC TRANS is >0.65 (the exact thresholds are up for debate). Then we are truly recording the state of the IFO and not relying on some script to write to the bit (I haven't thoguht through if there are some edge cases where we need an unreasonable number of diagnostic channels to determine if we are in a certain state or not).

Attachment 1: IFOSTATE.png
IFOSTATE.png
  16027   Wed Apr 14 13:16:20 2021 AnchalConfigurationComputers40m Control Room Changes
  • I have confirmed that the old two monitors' backlighting is not working. One can see the impression of the display without any brightness on them. Both old monitors are on the shelf behind.
  • Today we got a monitor and mouse from Mike. I had to change /etc/default/grub GRUB_GFXMODE to 1920x1200@30 on allegra for it to work with the(any) monitor.
  • Allegra is Debian 10 with latest cds-workstation installed on it. It is a good test station to migrate our existing scripts to start using updated cds-workstation configuration.
Quote:
  • Again, we have placed allegra's monitor for place holder but it is not working and we need new monitors for it in future whenever it is going to be used.

 

  16026   Wed Apr 14 13:12:13 2021 AnchalUpdateGeneralSorry, it was me

Sorry about that. It must be me. I'll make sure it doesn't happen again. I was careless to not check back, no further explanation.indecision

  16025   Wed Apr 14 12:27:10 2021 gautamUpdateGeneralLab left open again

Once again, I found the door to the outside in the control room open when I came in ~1215pm. I closed it.

  16024   Tue Apr 13 20:45:16 2021 YehonathanUpdatePSLLaser amplifier

{Yehonathan, Rana}

We unpacked the content of the amplifier crate in front of the water fountain (see attachments). Inside we found:

1. Amplifier head. (attachment 1)
2. Amplifier electronics and pump diodes (attachment 2).
3. Optical fiber (attachment 3).
4. 2 Long water hoses (~2m) and 2 short ones.
5. Network cable.
6. A wooden plate.
7. Cable sleeve (attachment 2)?
8. Some manuals will be uploaded to the wiki soon.

Please don't move/touch any of that stuff

Things that we need to consider/obtain:
1. A suitable power cable (attachment 4) with suitable power ratings (800W according to the amplifier specs). The connector head is C19 I believe.
2. A chiller. Rana says Aidan knows where to find one. Should we chill the amplifier head as well?
3. A mounting plate for the amplifier head with good thermal conductivity.
4. The pump wavelength is 808nm, we need to get suitable safety goggles.
5. Where to put the big electronics box. Preferably on 1X1 or 1X2.
6. How do we arrange the different components on the table? We also need to mode match the beam into the amplifier.

 

Attachment 1: 20210413_204721.jpg
20210413_204721.jpg
Attachment 2: 20210413_203300.jpg
20210413_203300.jpg
Attachment 3: 20210413_204940.jpg
20210413_204940.jpg
Attachment 4: 20210413_205549.jpg
20210413_205549.jpg
  16023   Tue Apr 13 19:24:45 2021 gautamUpdatePSLHigh power operations

We (rana, yehonathan and i) briefly talked about having high power going into the IFO. I worked on some calcs a couple of years ago, that are summarized here. There is some discussion in the linked page about how much power we even need. In summary, if we can have

  • T_PMC ~85% which is what I measured it to be back in 2019
  • T_IMC * T_inputFaraday ~60% which is what I estimate it to be now
  • 98% mode matching into the IMC
  • power recycling gain of 40-45 once we improve the folding mirror situation in the recycling cavities
  • and a gain of 270-280 in the arm cavities (20-30ppm round trip loss)

then we can have an overall gain of ~2400 from laser to each arm cavity (since the BS divides the power equally between the two arms). The easiest place to get some improvement is to improve T_IMC * T_inputFaraday. If we can get that up to ~90%, then we can have an overall gain of ~4000, which is I think the limit of what is possible with what we have.

We also talked about the EOM. At the same time, I had also looked into the damage threshold as well as clipping losses associated with the finite aperture of our EOM, which is a NewFocus 4064 (KTP is the Pockel medium). The results are summarized in Attachments #1 and #2 respectively. Rana thinks the EOM can handle factor of ~3 greater power than the rated damage threshold of 20W/mm^2.

Attachment 1: intensityDist.pdf
intensityDist.pdf
Attachment 2: clippingLoss.pdf
clippingLoss.pdf
  16022   Tue Apr 13 17:47:07 2021 gautamUpdateIOOWaveplate commissioning - software prepared

I spent some time today setting up a workable user interface to control the waveplate.

  1. Created some EPICS database records at /cvs/cds/caltech/target/ESP300.db. These are all soft channels. This required a couple of restarts of the modbus service on c1psl - as far as I can tell, everything has come back up without problems.
  2. Hacked newportESP to make it work, mainly some string encoding BS in the python2-->python3 paradigm shift.
  3. Made a python script at /cvs/cds/caltech/target/ESP300.py that is based on similar services I've set up for the CM servo and IMC servo boards. I have not yet set this up to run as a service on c1psl, but that is pretty trivial.
  4. Made a minimal MEDM screen, see Attachment #1. It is saved at  /opt/rtcds/caltech/c1/medm/c1psl/C1PSL_POW_CTRL.adl and can be accessed from the "PSL" tab on sitemap. We can eventually "calibrate" the angular position to power units.
  5. Confirmed that I can move the waveplate using this MEDM screen.

So this system is ready to be installed once Jordan and I find some time to lay out cabling + install the ESP300 controller in a rack.

At the moment, there is no high power and there is minimal risk of damaging anything, but someone should double check my logic to make sure that we aren't gonna burn the precious IFO optics. We should also probably hook up a hardware interlock to this controller.

I went through some aLIGO documentation and believe that they are using a custom made potentiometer based angle sensor rather than the integrated Newport (or similar) sensor+motor. My reading of the situation was that there were several problems to do with hysterisis, the "find home" routine etc. I guess for our purposes, none of these are real problems, as long as we are careful not to randomly rotate the waveplate through a full 180 degrees and go through the full fringe in the process. Need to think of a clever way to guard against careless / accidental MEDM button presses / slider drags.


Unrelated to this work: I haven't been in the lab for ~a week so I took the opportunity today to go through the various configs (POX/POY/PRMI resonant carrier etc). I didn't make a noise budget for each config but at least they can be locked 👍 . I also re-aligned the badly misaligned PMC and offloaded the somewhat large DC WFS offsets (~100 cts, which I estimate to be ~150 nNm of torque, corresponding to ~50 urad of misalignment) to the IMC suspensions' slow bias voltages. 

Attachment 1: remoteHWP.png
remoteHWP.png
  16021   Tue Apr 13 16:24:38 2021 Ian MacMillanUpdateCDS40m LSC simPlant model

Added Matlab to the Docker machine. This should help immensely with workflow as well as keeping installed libraries consistent. Next step is outlining the project so coding is easier

Command to launch is:     $ matlab &

From Jon just for bookkeeping: 

Then in the Matlab command window, open the CDS parts library via:

addpath /home/controls/simLink/lib/

open /home/controls/simLink/CDS_PARTS.mdl

Then open an RTCDS model (for example, here the LSC plant) via:

open /home/controls/docker-cymac/userapps/x1lsp.mdl

 

 

  16020   Tue Apr 13 09:51:22 2021 ranaUpdateSUSWhat's F2A??

Force to Angle. It just means the filters that are in the POS OUTPUT matrix. I think in the past sometimes they are called F2P or F2A.

These filters account for the frequency dependent coupling of the DOFs around the suspension resonance. Take a look at what Bhavini is doing for the plots.

 

  16019   Mon Apr 12 18:34:26 2021 YehonathanSummaryPSLPMC unlocked at 2pm on Sunday; ~ Restored

PMC lost lock again at around 16:00 April 12. I was able to lock it again but the transmission is only 0.6 now and REFL is 0.14.

Rana came in and realigned the PMC stirring mirrors. Now the transmission is 0.757V, and the REFL is 0.03V.

I noticed that the PZT was around 250V. Given that the PMC got unlocked at 16:00, which is around the peak temperature time in the lab (lagging behind the outside weather), due to the PZT voltage going down to 0V, I figured that the PZT voltage would go up during the night when the lab gets cold and therefore will likely go out of range again.

I found a different working point at 150V and relocked the PMC.

 

  16018   Mon Apr 12 17:30:11 2021 YehonathanUpdateBHDSOS assembly

Today, I screwed the plungers on the sensor plates and installed them on the Towers. I also installed the wire clamps on the suspension blocks (attachment).

I ran into problems in 2 separate suspension blocks: one had a dowel pin that was slightly too fat for the wire clamp. In another, the tapped holes were too short so that the 4-40 screws couldn't be screwed all the way.

 

Attachment 1: 20210412_170913.jpg
20210412_170913.jpg
  16017   Mon Apr 12 10:07:35 2021 AnchalUpdateSUSWhat's F2A??

I'm not sure I understand what F2A is? I couldn't find a description of this filter anywhere and don't remember if you have already explained it. Can you describe what is needed to be done again, please? We would keep SUS state space model and seismic transfer functions calculation ready meanwhile.

Quote:

Next we wanna get the F2A filters made since most of the IMC control happens at f < 3 Hz. Once you have the SUS state space model, you should be able to see how this can be done using only the free'swinging eigenfrequencies. Then you should get the closed loop model including the F2A filters and the damping filters to see what the closed loop behavior is like.

 

  16016   Mon Apr 12 08:32:54 2021 Anchal, PacoSummaryPSLPMC unlocked at 2pm on Sunday; ~ Restored

PMC lost lock between 21:00 and 22:00 UTC on April 11th as seen in the summary pages:

https://nodus.ligo.caltech.edu:30889/detcharsummary/day/20210411/psl/#gallery-4

That's between 2pm and 3pm on Sunday for us. We're not sure what caused it. We will attempt to lock it back.


Mon Apr 12 08:45:53 2021: we used milind's python script in scripts/PSL/PMC/pmc_autolocker.py. It locked the PMC in about a minute and then IMC catched lock succefully.

However, the PMC transmission PD shows voltage level of about 0.7V. On medm, it is set to turn red below 0.7 and yellow above. In Summary pages in the past, it seems like this value has typically been around 0.74V. Simil;arly, the reflection RFPD DC voltage is around 0.063 V right now while it is supposed to be around 0.04 nominally So the lock is not so healthy.

We tried running this script and the bashscript version too (scripts/PSL/PMC/PMCAutolocker) a couple of times but it was unable to acquire lock.

Then we manually tried to acquire lock by varying the C1:PSL-PMC_RAMP (with gain set to -10 dB) and resetting PZT position by toggling Blank. After a few attempts, we were able to find the lock with transmission PD value around 0.73V and reflection RFPD value around 0.043. PZT control voltage was 30V and shown in red in medm to begin with. So we adjusted the output ramp again to let it come to above 50V (or maybe it just drifted to that value by itself as we could se some slow drift too). At Mon Apr 12 09:50:12 2021 , the PZT voltage was around 58V and shown in green.

We assume this is a good enough point for PMC lock and move on.

  16015   Sat Apr 10 11:56:14 2021 JonUpdateCDS40m LSC simPlant model

Summary

Yesterday I resurrected the 40m's LSC simPlant model, c1lsp. It is running on c1sim, a virtual, self-contained cymac that Chris and I set up for developing sim models (see 15997). I think the next step towards an integrated IFO model is incorporating the suspension plants. I am going to hand development largely over to Ian at this point, with continued support from me and Chris.

x1lsp_main.png

 

LSC Plant

This model dates back to around 2012 and appears to have last been used in ~2015. According to the old CDS documentation:

Name Description Communicates directly with
LSP Simulated length sensing model of the physical plant, handles light propagation between mirrors, also handles alignment modeling and would have to communicate ground motion to all the suspensions for ASS to work LSC, XEP, YEP, VSP

Here XEP, YEP, and VSP are respectively the x-end, y-end, and vertex suspension plant models. I haven't found any evidence that these were ever fully implemented for the entire IFO. However, it looks like SUS plants were later implemented for a single arm cavity, at least, using two models named c1sup and c1spx (appear in more recent CDS documentation). These suspension plants could likely be updated and then copied for the other suspended optics.

To represent the optical transfer functions, the model loads a set of SOS filter coefficients generated by an Optickle model of the interferometer. The filter-generating code and instructions on how to use it are located here. In particular, it contains a Matlab script named opt40m.m which defines the interferferometer. It should be updated to match the parameters in the latest 40m Finesse model, C1_w_BHD.kat. The calibrations from Watts to sensor voltages will also need to be checked and likely updated.

Model-Porting Procedure

For future reference, below are the steps followed to port this model to the virtual cymac.

  1. Copy over model files.
     
    • The c1lsp model, chiara:/opt/rtcds/userapps/release/isc/c1/models/c1lsp.mdl, was copied to the userapps directory on the virtual cymac, c1sim:/home/controls/docker-cymac/userapps/x1lsp.mdl. In the filename, note the change in IFO prefix "c1" --> "x1," since this cymac is not part of the C1 CDS network.
       
    • This model also depends on a custom library file, chiara:/opt/rtcds/userapps/release/isc/c1/models/SIMPLANT.mdl, which was copied to c1sim:/home/controls/docker-cymac/userapps/lib/SIMPLANT.mdl.
       
  2. Update model parameters in Simulink. To edit models in Simulink, see the instructions here and also here.
     
    • The main changes are to the cdsParameters block, which was updated as shown below. Note the values of dcuid and specific_cpu are specifically assigned to x1lsp and will vary for other models. The other parameters will be the same.
       

    •  
    • I also had to change the name of one of the user-defined objects from "ADC0" --> "ADC" and then re-copy the cdsAdc object (shown above) from the CDS_PARTS.mdl library. At least in newer RCG code, the cdsAdc object must also be named "ADC0." This namespace collision was causing the compiler to fail.
       
    • Note: Since Matlab is not yet set up on c1sim, I actually made these edits on one of the 40m machines (chiara) prior to copying the model.
       
  3. Compile and launch the models. Execute the following commands on c1sim:
    • $ cd ~/docker-cymac
      $ ./kill_cymac
      $ ./start_cymac debug
    • The optional debug flag will print the full set of compilation messages to the terminal. If compilation fails, search the traceback for lines containing "ERROR" to determine what is causing the failure.
       

  4. Accessing MEDM screens. Once the model is running, a button should be added to the sitemap screen (located at c1sim:/home/controls/docker-cymac/userapps/medm/sitemap.adl) to access one or more screens specific to the newly added model.

    • Custom-made screens should be added to c1sim:/home/controls/docker-cymac/userapps/medm/x1lsp (where the final subdirectory is the name of the particular model).

    • The set of available auto-generated screens for the model can be viewed by entering the virtual environment:

    • $ cd ~/docker-cymac
      $ ./login_cymac                    #drops into virtual shell
      # cd /opt/rtcds/tst/x1/medm/x1lsp  #last subdirectory is model name
      # ls -l *.adl
      # exit                             #return to host shell
    • The sitemap screen and any subscreens can link to the auto-generated screens in the usual way (by pointing to their virtual /opt/rtcds path). Currently, for the virtual path resolution to work, an environment script has to be run prior to launching sitemap, which sets the location of a virtual MEDM server (this will be auto-scripted in the future):

    • $ cd ~/docker-cymac
      $ eval $(./env_cymac)
      $ sitemap
    • One important auto-generated screen that should be linked for every model is the CDS runtime diagnostics screen, which indicates the success/fail state of the model and all its dependencies. T1100625 details the meaning of all the various indicator lights.

 

 

 

 

  16014   Sat Apr 10 10:07:47 2021 ranaUpdateSUSFaster coil balancing

I think I mis-spoke about the balancing channels before. The ~20 Hz balancing could go into either the COIL banks or the SUS output matrix.

I believe its more conceptually clean to do this as gains in the outputmatrix, and leave the coil gains as +/- 1. i.e. we would only use the coil gains to compensate for coil/magnet actuation strength.

Then the high frequency balance goes into the outputmatrix. The F2A and A2L decoupling filters would then be generated having a high frequency gain = 1.

  16012   Sat Apr 10 08:51:32 2021 JonUpdateCDSI/O Chassis Assembly

I installed three of the 16-bit ADC adapter boards assembled by Koji. Now, the only missing hardware is the 18-bit DACs (quantities below). As I mentioned this week, there are 2-3 16-bit DACs available in the FE cabinet. They could be used if more 16-bit adapter boards could be procured.

C1BHD    
Component Qty Required Qty Installed
16-bit ADC 1 1
16-bit ADC adapter 1 1
18-bit DAC 1 0
18-bit DAC adapter 1 1
16-ch DIO 1 1
C1SUS2    
Component Qty required Qty Installed
16-bit ADC 2 2
16-bit ADC adapter 2 2
16-bit DAC 1 1
16-bit DAC adapter 1 1
18-bit DAC 5 0
18-bit DAC adapter 5 5
32-ch DO 6 6
16-ch DIO 1 1
  16011   Fri Apr 9 20:54:54 2021 YehonathanUpdateBHDSOS assembly

Today I assembled the skeleton of 6 towers, without clamps and sensor assembly (attachment 1).

Some of the side plates have this weird hole that doesn't fit any of the suspension blocks (attachment 2). I didn't notice when I counted the parts and now there are exactly enough side plates to assemble 7 towers.

Also found that one of the stiffener plates has a broken threading.

We will need more parts to go beyond the necessary 7 SOSs. I will do the recounting later.

Things to do next:

1. Find the capped spring plungers and send them to C&B.

2. Assemble the clamps onto the suspension blocks.

3. Push some Viton tips into the vented screws we got to make safety stops.

4. more C&B: Magnets, dumbells, dowel pins, OSEMs.

5. Push clean dowel pins into the last suspension block.

6. Assemble 7th Tower.

7. Assemble safety stops and clamps.

8. Glue magnets to dumbells.

 

 

Attachment 1: 20210409_202717.jpg
20210409_202717.jpg
Attachment 2: 20210409_202755.jpg
20210409_202755.jpg
  16010   Fri Apr 9 17:41:12 2021 ranaUpdateSUSFaster coil balancing

convergence is great.

Next we wanna get the F2A filters made since most of the IMC control happens at f < 3 Hz. Once you have the SUS state space model, you should be able to see how this can be done using only the free'swinging eigenfrequencies. Then you should get the closed loop model including the F2A filters and the damping filters to see what the closed loop behavior is like.

 

  16009   Fri Apr 9 13:13:00 2021 Anchal, PacoUpdateSUSFaster coil balancing

We ran again this method but with the 'b' parameter as a matrix instead. This provides more gain on some off-diagonal terms than others. This gave us a better convergence with the code reaching to the tolerance level provided (0.01 distance of S matrix from identity) within 16 iterations (~17 mins).


Attachment 1 again shows how the off-diagonal terms go down and how the overall distance of sensing matrix from identity goes down. This is 'Cross coupling budget' of the coils as iterations move forward.


Jumping to near zero-crossing:

  • Rana mentioned a ezlockin code which first makes 5 step changes in output matrix without using feedback and calculates the changes required to reach zero-crossing in the behavior of the off-diagonal terms during these steps.
  • This is similar to what we did above by hand where we increased the value of b for slowly converging off-diagonal elements.
  • We plan to implement this 'jump' to near zero-crossing method next. Aim is to get a coil balancing code that does the job in ~5 min.
  • We have been throwing away imaginary part of sensing matrix so far. We wanted to get to some owrking solution before we try more complex stuff. We have to figure out global phases in each transfer function estimate to rotate the measured transfer function appropriately.
Attachment 1: SmatIterations.pdf
SmatIterations.pdf
Attachment 2: MC2AllOutmat.txt
1.027604652272846142e+00 1.193175249772460367e+00 1.091939557371080394e+00
1.010054273887021292e+00 1.156057452309880551e+00 -8.392112351146234772e-01
9.895057930131009316e-01 -7.685799469766890768e-01 6.200896409311776880e-01
9.719554146272761930e-01 -8.056977444392685594e-01 -1.311061151554526294e+00
  16008   Thu Apr 8 20:58:17 2021 KojiUpdateCDSADC adapter boards assembly

5x 16bit ADC adapter boards (D0902006) assembled.

Attachment 1: P_20210408_205411_2_1.jpg
P_20210408_205411_2_1.jpg
  16007   Thu Apr 8 17:04:43 2021 Anchal, PacoUpdateSUSFirst Successful Coil Balancing

Today, we finally crossed the last hurdle and got a successful converging coil balancing run. laugh


What was the issue with POS?

  • Position of the MC2 mirror is being sensed using C1:IOO-MC_F_DQ channel which is proportional to the resonant frequency of the locked IMC.
  • However, this sensor is always 180 degrees out of phase of our actuator, the coils.
  • When the coils push the mirror forward, the length of the cavity actually decreases.
  • We added an extra option of providing a sign to the sensors such that -1 will be multiplied to sensed values for sensors which measure in opposite direction to the actuation.
  • This is important, because the feedback is applied to the coil output matrix assuming a particular direction of acctuation.
  • When we gave negative sign for the position sensor, it all started making sense and the algorithm started converging.

First run parameters:

  • We used binwidth of 0.25 Hz and duration of excitation as 41s. This would give welch and csd averaging of 19. We used median averaging to ignore outliers.
  • This iteration was run after PIT and YAW were separetly uncoupled before. We'll post a clean start to end run results in near future.
  • The iteration works in following manner:
    • Define a constant coil matrix C = [[1, 1, 1], [1, 1, -1], [1, -1, 1], [1, -1, -1]] which is ideal coil output matrix.
    • In each iteration, the output matrix Ok is defined as (note @ is the matmul operator):
      Ok = C @ Ak
      where Ak is a 3x3 matrix. A-1 is identity matrix.
    • At the end of each iteration, a sensing matrix is calculated in dimensions sensedDOF x excitedDOF, Sk
    • For next iteration, Ak+1 is calcualted by:
      Ak+1 = Ak - b * (Sk - I)
      where I is the identity matrix.
    • At convergence, the sensing matrix would become same as identity and matrix A will stop updating.
  • For this run, we kept the parameter b to be 0.05. This is similar to the KP parameter in PID loops. It should be between 0 and 1.
  • Since b value was small enough to allow for convergence from the inital point, but later it slowed down the process a lot.
  • Ideally, we should figure out a way to increase this paramter when the coil has been balanced somewhat, to increase the speed of the algorithm.
  • Secondly, we have a code which excites all DOFs at different frequencies directly using excitation channels in coil output matrix using awg.py. But for some reason, the excitation channel for 4th row in the output matrix column only connects intermittantly. Because of this, we can't use this method reliably yet. We can investigate more into it if suggested.

Balancing characteristics:

  • Attachment 1 shows how the distance of sensing matrix falls as iterations increase. We only ran for 50 iterations.
  • Attachment 2 shows how different off-diagonal terms of sensing matrix decreased.
    • Note that POS -> PIT, POS -> YAW and PIT-YAW have settled down to the noise floor.
    • The noise floor can be improved by increasing the excitation amplitude and/or increasing the duration of measurement.
  • Attachment 3 shows the evolution of sensing matrix as iterations move.

Final balanced output matrix:

Final balanced output coil matrix for MC2
POS PIT YAW COILS
1.02956 1.13053 1.19116 UL
1.01210 1.09188 -0.74832 UR
0.98737 -0.85502 0.70485 LR
0.96991 -0.89366 -1.23463 LR
Final Sensing Matrix
  Exc POS Exc PIT Exc YAW
Sens POS 1 -2.96e-2 8.00e-3
Sens PIT 8.58e-4 1 -4.84e-3
Sens YAW 5.97e-4 -1.15e-3 1

Code features and next:

  • Majority of the code is in two files: scripts/SUS/OutMatCalc/MC2crossCoupleTest.py and scripts/SUS/OutMatCalc/crossCoupleTest.py .
  • The code runs from start to end without human involevement and restores the state of channels in any case (error, kyboard interrupt, end of code) using finally statement.
  • Currently, each excitation is done one at a time through LockIn1. As mentioned above, this can be sped up 3 times if we get the awg.py to work reliably.
  • The complete code is in python3 and currently is run through native python3 on allegra (a new debian10 workstation with latest cds-workstation installed).
  • The code can be easily generalized for balancing any optic. Please let us know if we should work on making the generalized optic.
  • We're also working on thinking about increasing b as iterations move forward and the error signal becomes smaller.
  • We can also include the uncertainty in the Sensing matrix measurement to provide a weighted feedback. That way, we can probably increase b more.
Attachment 1: SDistanceFromIdentity.pdf
SDistanceFromIdentity.pdf
Attachment 2: SmatIterations.pdf
SmatIterations.pdf
Attachment 3: SmatrixPlots.pdf
SmatrixPlots.pdf SmatrixPlots.pdf SmatrixPlots.pdf SmatrixPlots.pdf SmatrixPlots.pdf SmatrixPlots.pdf SmatrixPlots.pdf SmatrixPlots.pdf
  16006   Wed Apr 7 22:48:48 2021 gautamUpdateIOOWaveplate commissioning

Summary:

I spent an hour today evening checking out the remote waveplate operation. Basic remote operation was established 👍 . To run a test on the main beam (or any beam for that matter), we need to lay out some long cabling, and install the controller in a rack. I will work with Jordan in the coming days to do these things. Apart from the hardware, some EPICS channel will need to be added to the c1ioo.db file and a python script will need to be set up as a service to allow remote operation.

Part numbers:

  • The controller is a NewFocus ESP300.
  • The waveplate stage is a PR50CC. The waveplate itself that is mounted has a 1" diameter (clear aperture is more like 21mm), which I think is ~twice the size of the waveplates we have in the lab, good thing Livingston shipped us the waveplate itself too. It is labelled QWPO-1064-10-2, so should be a half wave plate as we want, but I didn't explicitly check with a linearly polarized beam today. Before any serious high power tests, we can first contact clean the waveplate to avoid any burning of dirt. The damage threshold is rated as 1 MW/cm^2, and I estimate that we will be well below this threshold for any power levels (<30W) we are planning to put through this waveplate. For a 100um radius beam with 30W, the peak intensity is ~0.2 MW/cm^2. This is 20% of the rated damage threshold, so may be better to enforce that the beam be >200um going through this waveplate.
  • The dimensions of the mount look compatible with the space we have on the PSL table (though of course once the amplifier comes into the picture, we will have to change the layout. Maybe it's better to keep everything downstream of the PMC fixed - then we just re-position the seed beam (i.e. NPRO) and amplifier, and then mode-match the output of the amplifier to the PMC.

Electrical tests:

  1. First, I connected a power cord to the ESP300 and powered it on - the front display lit up and displayed a bunch of diagnostics, and said something to the effect of "No stage connected".
  2. Next, I connected the rotary mount to "Axis #1": Male DB25 on the stage to female DB25 on the rear of the ESP300. The stage was recognized.
  3. Used the buttons on the front panel to rotate the waveplate, and confirmed visually that rotation was happening 👍 . I didn't calibrate the actual degrees of rotation against the readback on the front panel, but 45 degrees on the panel looked like 45 degrees rotation of the physical stage so seems fine.

RS232 tests:

  • This unit only has a 9-pin Dsub connector to interface remotely to it, via RS232 protocol. c1psl Supermicro host was designated the computer with which I would attempt remote control.
  • To test, I decided to use a serial-USB adapter. Since this is only a single unit, no need to get an RS232-ethernet interface like the one used in the vacuum rack, but if there are strong opinions otherwise we can adopt some other wiring/control philosophy.
  • No drivers needed to be installed, the host recognized the adapter immediately. I then shifted the waveplate and controller assembly to inside the VEA - they are sitting on a cart behind 1X2. Once the controller was connected to the USB-serial adapter cable, it was registered at /dev/ttyUSB0 immediately. I had to chown this port to the controls user for accessing it using python serial
  • Initially, I was pleasantly surprised when I found not one but TWO projects on PyPi that already claimed to do what I want! Sadly, neither NewportESP1.1 nor PyMeasure0.9.0 actually worked - the former is for python2 (and the string typesetting has changed for PySerial compatible with python3), while the latter seems to be optimized for Labview interfacing and didn't play so nice with the serial-USB adapter. I didn't want to spend >10mins on this and I know enough python serial to do the interfacing myself, so I pushed ahead. Good thing we have several pySerial experts in the group now, if any of you want to figure out how we can make either of these two utilities actually work for us - there is also this repo which claims to work for python 3 but I didn't try it because it isn't a managed package.
  • The command list is rather intimidating, it runs for some 100 (!) pages. Nevertheless, I used some basic commands to readback the serial number of the controller, and also succeeded in moving the stage around  by issuing the "PR" command appropriately 👍. BTW, I forgot that I didn't test the motor enable/disable which is an essential channel I think.
  • I think we actually only need a very minimal set of commands, so we don't need to read all 100 pages of instructions:
    • motor enable/disable
    • absolute and relative rotations
    • readback of the current position
    • readback of the moving status
    • a stop command
    • an interlock
  • Note that as a part of this work, in addition to chowning /dev/ttyUSB0, I installed the two aforementioned python packages on c1psl. I saw no reason to manually restart the modbus and latch services running on it, and I don't believe this work would have impacted the correct functioning of either of those two services, but be aware that I was poking around on c1psl. I was also reminded that the system python on this machine is 2.7 - basically, only the latch service that takes care of the gains for the IMC servo board are dependent on python (and my proposed waveplate control script will be too), but we should really upgrade the default python to 3.7/3.8.

Next steps:

Satisfied that the unit works basically as expected, I decided to stop for today. My thinking was that we can have the ESP300 installed in 1X1 or 1X2 (depending on where space is more readily available). I will upload have uploaded a cartoon here so people can comment if they like/dislike my plan

  • We need to use a long-ish cable to run from 1X1/1X2, where the controller will be housed, to the PSL enclosure. Livingston did ship one such long cable (still on Rana's table), but I didn't check if the length is sufficient / the functionality of this long cable. 
  • We need to set up some EPICS channels for the rotation stage angle, motor ENABLE/DISABLE, a "move stage" button, motion status, and maybe a channel to control the rotation speed? 
  • We need a python script that is reading from / writing to these EPICS channel in a while loop. Should be straightforward to setup something to run like the latch.py service that has worked decently reliably for ~a year now. afaik, there isn't a good way to run this synchronously, and the delay in sending/completing the execution of some of the serial commands might be ~1 second, but for the purpose of slowly ramping up the power, this shouldn't be a problem.
  • One question I do have is, what is the strategy to protect the IFO from the high power when the lock is lost? Surely we are not gonna rely on this waveplate for any fast actuation? With the current input power of 1W, the MCREFL photodiode sees ~100mW when the IMC loses lock. So if the final input power is 35W, do we wanna change the T=10% beamsplitter in the MCREFL path to keep this ratio?

Once everything is installed, we can run some tests to see if the rotary motion disturbs the PSL in any meaningful way. I will upload some photos to the picasa later. Photos here.

Attachment 1: remotePowCtrl.pdf
remotePowCtrl.pdf
  16005   Wed Apr 7 17:38:51 2021 AnchalUpdateSUSTrying to uncouple only PIT and YAW first

To test if our method is working at all, we went for the simpler case of just uncoupling PIT and YAW. This is also because the sensor used for these two degrees of freedom is similar (the MC Trans WFS).

We saw a successful decrease in cross-coupling between PIT and YAW over the first 50 iterations that we tried. Here are some results:


Final output matrix:

Output matrix for uncoupling PIT and YAW from eachother
PIT YAW COILS
1.01858 1.16820 UL
0.98107 -0.79706 UR
-0.98107 0.79706 LL
-1.01858 -1.16820 LR

Plots:

  • Attachment 1 shows distance of sensing matrix from identity as iterations go.
  • Attachment 2 shows the off-diagonal elements of sensing matrix as the iterations increase.
    • It is worth noting that PIT -> YAW coupling was the main element that was reduced successfully while the YAW -> PIT was reducing but much more slowly.
    • Most of the remaining cross coupling in the end was from YAW -> PIT.
  • Attachment 3 shows first 10 oscillations in the time series data during excitation of some of the iterations.
  • Attachment 4 shows the cross spectral density of the sensed data during excitation with each other. This has been normalized by reference PSD data (taken with no excitation) of the sensed DOFs involved in the CSD calculation.
  • Attachment 5 shows the TF estimate made by normalizing CSD data column wise by the diagonal elements. The excitation frequency point in these plots become the Sensing matrix in the calculation.
    • One can notice how the PIT -> YAW element is going down in these plots.
    • Even though we are using only the real value of the sensing matrix, the imaginary values are also going down.

Next, tried uncoupling POS and PIT:

  • Next, we tried to uncouple POS and PIT. We expect them to be more coupled than with YAW.
  • At the time of writing this post, 15 iterations of this attempt have been completed and it is not looking good sad.
  • The distance of the sensing matrix from identity is growing at an accelerated rate.
  • The POS output matrix column seems to be trying to go towards the negative of PIT output matrix column! Why? We don't know.
  • We have seen in the past that once POS transforms into PIT or YAW, it just makes the output matrix worse as no feedback actually goes into the POS column. Eventually, the IMC will cease to remain locked.
  • So, I'm cancelling this attempt for now. Will consider more alternatives later.
Attachment 1: SDistanceFromIdentity.pdf
SDistanceFromIdentity.pdf
Attachment 2: SmatIterations.pdf
SmatIterations.pdf
Attachment 3: TimeSeriesPlots.pdf
TimeSeriesPlots.pdf TimeSeriesPlots.pdf TimeSeriesPlots.pdf TimeSeriesPlots.pdf TimeSeriesPlots.pdf TimeSeriesPlots.pdf TimeSeriesPlots.pdf TimeSeriesPlots.pdf
Attachment 4: CSDPlots.pdf
CSDPlots.pdf CSDPlots.pdf CSDPlots.pdf CSDPlots.pdf CSDPlots.pdf CSDPlots.pdf CSDPlots.pdf CSDPlots.pdf
Attachment 5: SmatrixPlots.pdf
SmatrixPlots.pdf SmatrixPlots.pdf SmatrixPlots.pdf SmatrixPlots.pdf SmatrixPlots.pdf SmatrixPlots.pdf SmatrixPlots.pdf SmatrixPlots.pdf
  16004   Wed Apr 7 13:07:03 2021 JordanUpdateSUSCoM on 3"->2" Adapter Ring for SOS

Adding the chamfer around the edge of the optic ring did not change the center of mass relative to the plane from the suspension wires.

The CoM was .0003" away from the plane. Adding the chamfer moved it closer by .0001". See the attached photo.

I've also attached the list of the Moments of Inertia of the SOS Assembly.

Attachment 1: CoM_with_Chamfer.png
CoM_with_Chamfer.png
Attachment 2: Moments_of_Inertia.PNG
Moments_of_Inertia.PNG
  16003   Wed Apr 7 02:50:49 2021 KojiUpdateSUSFlange Inspections

Basically I went around all the chambers and all the DB25 flanges to check the invac cable configurations. Also took more time to check the coil Rs and Ls.

Exceptions are the TTs. To avoid unexpected misalignment of the TTs, I didn't try to disconnect the TT cables from the flanges.

Upon the disconnection of the SOS cables, the following steps are taken to avoid large impact to the SOSs

  • The alignment biases were saved or recorded.
  • Gradually moved the biases to 0
  • Turned off the watchdogs (thus damping)

After the measurement, IMC was lock and aligned. The two arms were locked and aligned with ASS. And the PRM alignment (when "misalign" was disengaged) was checked with the REFL CCD.
So I believe the SOSs are functioning as before, but if you find anything, please let me know.
 

  16002   Tue Apr 6 21:17:04 2021 KojiSummaryGeneralPSL HEPA investigation

- Last week we found both of the PSL HEPA units were not running.

- I replaced the capacitor of the north unit, but it did not solve the issue. (Note: I reverted the cap back later)
- It was found that the fans ran if the variac was removed from the chain.
- But I'm not certain that we can run the fans in this configuration with no attendance considering fire hazard.

@3AM: UPON LEAVING the lab, I turned off the HEPA. The AC cable was not warm, so it's probably OK, but we should wait for the continuous operation until we replace the scorched AC cable.


The capacitor replacement was not successful. So, the voltages on the fan were checked more carefully. The fan has the three switch states (HIGH/OFF/LOW). If there is no load (SW: OFF), the variac out was as expected. When the load was LOW or HIGH, it looked as if the motor is shorted (i.e. no voltage difference between two wires).

I thought the motors may have been shorted. But if the load resistance was measured with the fluke meter, it showed some resistance

- North Unit: SW LOW 4.6Ohm / HIGH 6.0Ohm
- South Unit: SW LOW 6.0Ohm / HIGH 4.6Ohm (I believe the internal connection is incorrect here)

I believed the motors are alive! Then the fans were switched on with the variac removed... they ran. So I set the switch LOW for the north unit and HIGH for the south unit.

Then I inspected the variac:

  • The AC output has some liquid leaking (oil?) (Attachment 1)
  • The AC plug on the variac out has a scorch mark (Attachments 2/3)

So, this scorched AC plug/cable connected directly to the AC right now. I'm not 100% confident about the safety of this configuration.
Also I am not sure what was wrong with the system.

  • Has the variac failed first? Because of the heat? I believe that the HEPA was running @30% most of the time. Maybe the damage was already there at the failure in Nov 2020?
  • Or has the motor stopped at some point and this made the variac failed?
  • Was the cable bad and the heat made the variac failed (then the problem is still there).

So, while I'm in the lab today, I'll keep the HEPA running, but upon my taking off, I'll turn it off. We'll discuss what to do in the meeting tomorrow.

 

Attachment 1: 20210406211741_IMG_0554.jpeg
20210406211741_IMG_0554.jpeg
Attachment 2: 20210406211840_IMG_0555.jpeg
20210406211840_IMG_0555.jpeg
Attachment 3: 20210406211850_IMG_0556.jpeg
20210406211850_IMG_0556.jpeg
ELOG V3.1.3-